# Interplay between Majorana and Shiba states in a minimal Kitaev chain coupled to a superconductor

### v1

@@ Appendix A Poles Structure and EPs
Figure S 2 illustrates the associated physics when detuning QD2. The excited states obeys that when $\Gamma_{s}/t\rightarrow\infty$ then $z_{s,e}\rightarrow\varepsilon_{2}$. More in particular, when $\varepsilon_{2}<\Delta$, the excited state is defined by a renormalized YSR state that do not merge with the MZMs, see Fig. S 2(a). Fig. S 2(d,e) shows that in the ultra strong coupling regime, the excited subgap state becomes strongly localized at QD2. Moreover, there is no longer secondary EPs for large couplings, which can be related to the lack of a quantized transport response, leading to a vanishing normal transport signal in the ultra-strong coupling regime (not shown). By contrast, when $\varepsilon_{2}\geq\Delta$, the excited state becomes pinned to $z_{s,e}=\varepsilon_{2}$ even at the strong coupling transition $\Gamma_{s}/t>1$.
@@
Analyzing the dispersion of the excited state with $\varepsilon_{2}$,, for moderate couplings the excited state follows a parabolic dispersion, Fig. S 2(b). However, as the coupling with the SC lead is increased, the minimum of the parabola is displaced to smaller energy values and the dispersion is slightly renormalized. In the limit $\Gamma_{s}\rightarrow\infty$, the excited state becomes linear with $\varepsilon_{2}$, and the energy minimum finally collapses to zero, see Fig. S 2(c). By contrast, when detuning QD1, the parabolic dispersion of the excited state with $\varepsilon_{1}$ becomes constant, as QD1 is being absorbed by the SC lead for large couplings (not shown).
@@
Figure S 3: (a) Real and imaginary part (straight and dotted lines respectively) of the poles as a function of the hybridization parameter to the SC lead obtained when considering a finite polarization in QD1. (b-d) EPs for $\lambda_{s}/t=5,10,15$ marked by arrows in (a). (e-g) LDOS across the junction, showing the spreading of the states into the SC lead. White dashed lines represents the pole obtained computationally from the characteristic polynomial in Eq. (10) when considering a finite polarization. We set $\lambda_{n}/t=1$ and $E_{z}/t=10$ in all panels.
@@ Appendix B Penetration of subgap states
We follow Ref. 48 to analyze the penetration of the subgap poles into the leads from the total advanced GF to compute the Local Density of States (LDOS) as $\rho(\omega)=\operatorname{\mathfrak{Im}}\hat{G}_{\bar{j}\bar{j}}/\pi$ in the $4\times 4$ Nambu basis, where the GF in the semi-infinite SC lead acquires the form
$\hat{G}_{\bar{j}\bar{j}}=\hat{{\cal G}}_{\bar{j}\bar{j}}+\hat{{\cal G}}_{\bar{j}0}\hat{\Sigma}_{s}[\hat{\mathbb{I}}-\hat{{\cal G}}_{11}\hat{\Sigma}_{s}^{\dagger}\hat{{\cal G}}_{s}\hat{\Sigma}_{s}]^{-1}\hat{{\cal G}}_{11}\hat{\Sigma}_{s}^{\dagger}\hat{{\cal G}}_{0\bar{j}},$ (14)
where $\bar{j}=-j$ is an integer position index running into the left SC lead and $\hat{T}_{s}=t_{s}\sigma_{z}\tau_{0}$ being the coupling between neighbouring sites inside the SC lead. The GFs associated to the isolated semi-infinite SC lead takes the expression
$\hat{{\cal G}}_{\bar{j}\bar{j}}=$ $\,\hat{G}^{(0)}_{00}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{G}^{(0)}_{\bar{j}0}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{0\bar{j}}\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}\ ,$ (15a)
$\hat{{\cal G}}_{\bar{j}0}=$ $\,\hat{G}^{(0)}_{\bar{j}0}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{G}^{(0)}_{\bar{j}0}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{10},$ (15b)
$\hat{{\cal G}}_{0\bar{j}}=$ $\,\hat{G}^{(0)}_{0\bar{j}}-\hat{G}^{(0)}_{01}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{0\bar{j}}\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}\ .$ (15c)
@@
$\hat{G}^{(0)}_{00}=$ $\,[\hat{\mathbb{I}}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{{\cal G}}_{s}\hat{T}_{s}^{\dagger}]^{-1}\hat{{\cal G}}_{s}\ ,$ (16a)
$\hat{G}^{(0)}_{\bar{j}0}=$ $\,[\hat{{\cal G}}_{s}\hat{T}_{s}]^{j}\hat{G}^{(0)}_{00},$ (16b)
$\hat{G}^{(0)}_{0\bar{j}}=$ $\,\hat{G}^{(0)}_{00}[\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}]^{j}.$ (16c)
It should be noticed that to obtain a proper convergence of the GFs across the SC lead, $\hat{{\cal G}}_{s}$ is defined out of the wide-band approximation, i.e., the normal and anomalous part of the bGF are
$g=$ $\,\frac{\omega}{2t_{s}^{2}}\left(1-\sqrt{\frac{4t_{s}^{2}+\Delta_{s}^{2}-\omega^{2}}{\Delta_{s}^{2}-\omega^{2}}}\right),$ (17a)
$f=$ $\,\frac{\Delta_{s}}{2t_{s}^{2}}\left(1-\sqrt{\frac{4t_{s}^{2}+\Delta_{s}^{2}-\omega^{2}}{\Delta_{s}^{2}-\omega^{2}}}\right).$ (17b)
It is trivial to see that, to compute the penetration into the normal lead, one has to transform the previous expressions following $\bar{j}\rightarrow j$ where $\hat{\Sigma}_{s}\rightarrow\hat{\Sigma}_{n}^{\dagger}$, $\hat{{\cal G}}_{11}\rightarrow\hat{{\cal G}}_{22}$ , $\hat{{\cal G}}_{s}\rightarrow\hat{{\cal G}}_{n}$ and $\hat{T}_{s}\rightarrow\hat{T}_{n}^{\dagger}=t^{*}_{n}\sigma_{z}\tau_{0}$. The normal bGF can be obtained from Eq. (17a) considering $\Delta_{s}=0$, where $\hat{{\cal G}}_{n}=(x+i\sqrt{1-x^{2}})/t_{n}\tau_{0}$, and $x=\omega/(2t_{n})$ for advanced GFs.
@@ Appendix C Finite spin-polarization on QD1
In the main text we have discussed a configuration where QD1 acquires a finite spin polarization, leading to the coexistence of the YSR and MZMs occurring at much smaller couplings with the SC lead. After the YSR minima of energy, a destructive interference occurs at $\Gamma_{s}\propto E_{z}$, after which the quantized zero bias peak in the normal conductance disappears. To analyze this configuration we have to consider that there are different Zeeman fields for each QD in Eq. (4), where $E_{z,1}=E_{z}$ is finite and, $E_{z,2}\rightarrow\infty$. Notice that the GF describing QD1 has to be described in the full $4\times 4$ Nambu structure.
Figure S 3 shows the poles of the system with partial polarization for different values of the coupling to the SC lead. In the ultra strong coupling regime a new transition at $\Gamma_{s}\sim E_{z}$ occurs, where a pair of poles detach from zero and the bifurcation of imaginary parts opened at $\Gamma_{s}/t\sim 4$, is finally closed for negative values $\operatorname{\mathfrak{Im}}z<0$, see Fig. S 3(a). This new regime induces a destructive interference between the YSR and the MZMs as there are no longer secondary EPs for $\operatorname{\mathfrak{Im}}z<0$ above the transition, see Fig. S 3(c). Moreover, non-zero values of $\Gamma_{-}>0$ for the main EP, and a strong renormalizacion of the positive valued secondary EP $\operatorname{\mathfrak{Im}}z>0$ is observed in Fig. S 3(d), signalling the spoiling of the Majorana properties after the crossing.
Concerning the degree of localization of the subgap states across the junction, Fig.S 3(e) shows the coexistence of the YSR and the MZMs for small couplings, where the Majorana living in QD1 is strongly delocalized into the SC lead. At the maximum approach of both subgap states $\Gamma_{s}\sim E_{z}$ in Fig. S 3(f), the YSR loses the oscillatory behaviour. For even larger values of the coupling, one finds the destructive interference regime where the MZMs living in QD1 disappears, giving as a result a trivial zero pole in QD2 and a YSR state.
@@ Appendix D Superconducting Transport
For computing transport properties in junctions, we start from the general expression for the mean value of the current using Keldysh Green functions in Refs. 46, 31. When considering the current through different SCs, it is convenient to define a Gauge in which the chemical potential difference appears as a time-dependent tunnel coupling, where the SC phase difference evolves as $\phi(t)=\phi_{0}+(2eV/\hbar)t$, cf. Ref. 31. By a suitable choice of basis, the coupling between the SC probe and a fully polarized Kitaev chain only involves one spin species, equivalent to remove the anomalous part of the SC lead as no multiple Andreev reflections are allowed [51, 30, 31] (i.e., the probe act as a normal lead with a gap). Besides, there is still resonant Andreev transport in the minimal Kitaev chain due to p-wave correlations. Thus, the current through a BCS probe and a spinless p-wave SC acquires the form,
@@
$G_{RR}(\omega)={}$ $\left[[{\cal G}_{R}]^{-1}-\Sigma(\omega)\right]^{-1},$ (19a)
$\Sigma(\omega)={}$ $\Sigma_{LR}^{\dagger}\bar{{\cal G}_{s}}(\omega)\Sigma_{LR},$ (19b)
@@
$\bar{{\cal G}_{s}}(\omega)=\operatorname{diag}\left(\frac{-\omega_{-}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega_{-}^{2}}},\frac{-\omega_{+}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega_{+}^{2}}}\right),$ (20)
@@
$G^{K}_{RR}(\omega)=G^{\dagger}_{RR}\left\{\Sigma^{K}+[{\cal G}^{\dagger}_{R}]^{-1}{\cal G}^{K}_{R}\ [{\cal G}_{R}]^{-1}\right\}G_{RR},$ (21)
@@
${\cal G}^{K}_{R}(\omega)={}$ $(1-2f(\omega))\big{(}{\cal G}^{\dagger}_{R}-{\cal G}_{R}\big{)},$ (22a)
$\Sigma^{K}(\omega)={}$ $\Sigma_{LR}^{\dagger}F_{\omega}\big{(}\bar{{\cal G}_{s}}^{\dagger}-\bar{{\cal G}_{s}}\big{)}\Sigma_{LR},$ (22b)
@@
$\tilde{\Sigma}^{R/A}(\omega)=\Sigma^{R/A}_{ee}(\omega)\Pi_{e\uparrow},\quad\tilde{\Sigma}^{K}(\omega)=\Sigma^{K}_{ee}(\omega)\Pi_{e\uparrow},$ (23)
where we have defined the projector $\Pi_{e\uparrow}=(\sigma_{0}+\sigma_{z})/2$, as only up-spin electrons from the BCS lead are tunnel coupled to the spinless fermions from the minimal Kitaev chain. Finally, we extract the differential conductance from the current using that $G_{S}=\partial I_{S}/\partial V$ in the $T\rightarrow 0$ limit.
The small but finite temperature in the calculation allows for thermally excited quasiparticles from the continuum, inducing a broadening term ($\Gamma_{sp}$) of the Majorana zero pole and enabling single particle current by means of inelastic cotunneling. Furthermore, there is an extra relaxation contribution coming from the non-local coupling to the normal lead. It should be noticed that, for a zero pole, both resonant Andreev (i.e., virtual occupation of the state) and single particle processes (i.e., occupation of the state requiring relaxation to empty the state) share the same threshold signal at $e|V|=\Delta$. Notice that, along with the inelastic cotunneling, the hybridization with the SC lead induces a voltage bias dependent broadening ($\Gamma_{e/h}$) of the Majorana zero mode, which may modify the position of the conductance peak related to the threshold. The dominant contribution to the broadening of the subgap state defines two different regimes [30, 53]:

### v3

@@ Appendix A Poles Structure and EPs
Figure S 2 illustrates the strongly renormalized physics when detuning QD2. The excited states obey that when $\Gamma_{s}/t\rightarrow\infty$ then $z_{s,e}\rightarrow\varepsilon_{2}$. More in particular, when $\varepsilon_{2}<\Delta$, the excited state is defined by a renormalized YSR state that do not merge with the MZMs, see Fig. S 2(a). Fig. S 2(d,e) shows that in the ultra strong coupling regime, the excited subgap state becomes strongly localized at QD2. Moreover, there is no longer secondary EP for large couplings (not shown), which can be related to the lack of a quantized transport response, leading to a vanishing normal transport signal in the ultra-strong coupling regime, cf. App. G. By contrast, when $\varepsilon_{2}\geq\Delta$, the excited state becomes pinned to $z_{s,e}=\varepsilon_{2}$ even at the strong coupling transition $\Gamma_{s}/t>1$.
@@
Analyzing the dispersion of the excited state with $\varepsilon_{2}$, for moderate couplings the excited state follows a parabolic dispersion, Fig. S 2(b). However, as the coupling with the SC lead is increased, the minimum of the parabola is displaced to smaller energy values and the dispersion is slightly renormalized. In the limit $\Gamma_{s}\rightarrow\infty$, the excited state becomes linear with $\varepsilon_{2}$, and the energy minimum finally collapses to zero, see Fig. S 2(c). By contrast, when detuning QD1, the parabolic dispersion of the excited state with $\varepsilon_{1}$ becomes constant, as QD1 is being absorbed by the SC lead for large couplings (not shown).
@@
Figure S 2: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead going to the ultra-strong coupling for $\varepsilon_{2}/t=0.5$. (b-c) LDOS as a function of QD2 level for $\lambda_{s}/t=10,40$ marked by arrows in (a). (d-e) LDOS across the junction, showing the spreading of the states into the SC lead. White dashed lines represents the pole obtained computationally from the characteristic polynomial in Eq. (S 9). We set $\lambda_{n}/t=1$ in all panels.
@@ Appendix B Penetration of Subgap States
We follow Ref. [66] to analyze the penetration of the subgap poles into the leads from the total advanced GF to compute the Local Density of States (LDOS) as $\rho(\omega)=\operatorname{\mathfrak{Im}}\hat{G}_{\bar{j}\bar{j}}/\pi$ in the $4\times 4$ Nambu basis (hat notation), where the GF in the semi-infinite SC lead acquires the form
$\hat{G}_{\bar{j}\bar{j}}=\hat{{\cal G}}_{\bar{j}\bar{j}}+\hat{{\cal G}}_{\bar{j}0}\hat{\Sigma}_{s}[\hat{\mathbb{I}}-\hat{{\cal G}}_{11}\hat{\Sigma}_{s}^{\dagger}\hat{{\cal G}}_{s}\hat{\Sigma}_{s}]^{-1}\hat{{\cal G}}_{11}\hat{\Sigma}_{s}^{\dagger}\hat{{\cal G}}_{0\bar{j}},$ (S 13)
where $\bar{j}=-j$ is an integer position index running into the left SC lead and $\hat{T}_{s}=t_{s}\sigma_{z}\tau_{0}$ being the coupling between neighbouring sites inside the SC lead. The bGFs associated to the isolated semi-infinite SC lead takes the expression
$\hat{{\cal G}}_{\bar{j}\bar{j}}=$ $\,\hat{G}^{(0)}_{00}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{G}^{(0)}_{\bar{j}0}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{0\bar{j}}\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}\ ,$ (S 14a)
$\hat{{\cal G}}_{\bar{j}0}=$ $\,\hat{G}^{(0)}_{\bar{j}0}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{G}^{(0)}_{\bar{j}0}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{10},$ (S 14b)
$\hat{{\cal G}}_{0\bar{j}}=$ $\,\hat{G}^{(0)}_{0\bar{j}}-\hat{G}^{(0)}_{01}\big{[}\hat{G}^{(0)}_{00}\big{]}^{-1}\hat{G}^{(0)}_{0\bar{j}}\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}\,.$ (S 14c)
@@
$\hat{G}^{(0)}_{00}=$ $\,[\hat{\mathbb{I}}-\hat{{\cal G}}_{s}\hat{T}_{s}\hat{{\cal G}}_{s}\hat{T}_{s}^{\dagger}]^{-1}\hat{{\cal G}}_{s}\ ,$ (S 15a)
$\hat{G}^{(0)}_{\bar{j}0}=$ $\,[\hat{{\cal G}}_{s}\hat{T}_{s}]^{j}\hat{G}^{(0)}_{00}\,,$ (S 15b)
$\hat{G}^{(0)}_{0\bar{j}}=$ $\,\hat{G}^{(0)}_{00}[\hat{T}_{s}^{\dagger}\hat{{\cal G}}_{s}]^{j}\,.$ (S 15c)
It should be noticed that to obtain a proper convergence of the GFs across the SC lead, the bGF of the latter is defined out of the wide-band approximation, i.e., the normal and anomalous part of the bGF, $\hat{{\cal G}}_{s}=g(\omega)\,\sigma_{0}\tau_{0}-f(\omega)\,\sigma_{y}\tau_{y}$, are
$g(\omega)=$ $\,\frac{\omega}{2t_{s}^{2}}\left(1-\sqrt{\frac{4t_{s}^{2}+\Delta_{s}^{2}-\omega^{2}}{\Delta_{s}^{2}-\omega^{2}}}\right),$ (S 16a)
$f(\omega)=$ $\,\frac{\Delta_{s}}{2t_{s}^{2}}\left(1-\sqrt{\frac{4t_{s}^{2}+\Delta_{s}^{2}-\omega^{2}}{\Delta_{s}^{2}-\omega^{2}}}\right).$ (S 16b)
It is trivial to see that, to compute the penetration into the normal lead, one has to transform the previous expressions following $\bar{j}\rightarrow j$ where $\hat{\Sigma}_{s}\rightarrow\hat{\Sigma}_{n}^{\dagger}$, $\hat{{\cal G}}_{11}\rightarrow\hat{{\cal G}}_{22}$ , $\hat{{\cal G}}_{s}\rightarrow\hat{{\cal G}}_{n}$ and $\hat{T}_{s}\rightarrow\hat{T}_{n}^{\dagger}=t^{*}_{n}\sigma_{z}\tau_{0}$. The normal bGF can be obtained from Eq. (S 16a) considering $\Delta_{s}=0$, where $\hat{{\cal G}}_{n}=(x+i\sqrt{1-x^{2}})/t_{n}\,\sigma_{0}\tau_{0}$, and $x=\omega/(2t_{n})$. Notice that the analyticity of the advanced GFs imposes a change of sign of the square root for bGF of the normal lead with respect to the SC case.
@@ Appendix C Finite Spin-Polarization on QD1
Figure S 3: (a) Real and imaginary part (solid and dotted lines respectively) of the poles as a function of the hybridization parameter to the SC lead obtained when considering a finite polarization in QD1. (b-d) EPs for $\lambda_{s}/t=5,10,15$ marked by arrows in (a). (e-g) LDOS across the junction, showing the spreading of the states into the SC lead. White dashed lines represents the pole obtained computationally from the characteristic polynomial in Eq. (S 9) when considering a finite polarization. Inset shows a close-up image around $\Gamma_{s}\sim E_{z}$ of the imaginary part of the poles. We set $\lambda_{n}/t=1$ and $E_{z}/t=10$ in all panels, except indicated, we have set $t/\Delta=1$.
In the main text, we have discussed a configuration where QD1 acquires a finite spin polarization, leading to the coexistence of the YSR and MZMs occurring at much smaller couplings with the SC lead. After the YSR minima of energy, a destructive interference occurs at $\Gamma_{s}\propto E_{z}$, after which the quantized zero-bias peak in the normal conductance disappears. To analyze this configuration, we have to consider that there are different Zeeman fields for each QD in Eq. (S 3), where $E_{z,1}=E_{z}$ is finite and, $E_{z,2}\rightarrow\infty$. Notice that the GF describing QD1 has to be described in the full $4\times 4$ Nambu structure.
Figure S 3 shows the poles of the system with partial polarization for different values of the coupling to the SC lead. We focus now on the negative part of the poles as it appears a new transition at $\Gamma_{s}\sim E_{z}$, where a bifurcation of imaginary parts closes and reopens, see Fig. S 3(a). Before the transition, the system shows a similar EPs structure as in the fully polarized case where the maximum (minimum) value of the secondary bifurcation is $\Gamma^{+}_{2}\approxeq\Gamma_{s}$ ($\Gamma^{-}_{2}=\Gamma_{n}$), see Fig. S 3(b). However, it shows non-zero values of $\Gamma^{-}_{1}\gtrapprox 0$ for the main EP. This distorted EP structure for the finite spin polarization case could explain the lack of YSR quantized conductance signal for this regime in contrast to the fully polarized case.
At the transition, the new regime induces a destructive interference between the YSR and the MZMs, see Fig. S 3(c). Moreover, the non-zero values of $\Gamma^{-}_{1}>0$ for the main EP increases and, for the secondary EP, $\Gamma^{+}_{2}$ decreases, satisfying that $\Gamma_{s}\gg\Gamma^{+}_{2}>\Gamma_{n}$. Above the transition, Fig. S 3(d) shows a strong renormalization of the EPs such $\Gamma^{+}_{2},\,\Gamma^{-}_{1}\rightarrow\Gamma_{n}$ as both of them are collapsing to the same point, signalling the spoiling of the Majorana properties after the crossing point.
Concerning the degree of localization of the subgap states across the junction, Fig.S 3(e) shows the coexistence of the YSR and the MZMs for small couplings, where the Majorana living in QD1 is strongly delocalized into the SC lead. At the maximum approach of both subgap states $\Gamma_{s}\sim E_{z}$ in Fig. S 3(f), the YSR loses the oscillatory behaviour [76]. For even larger values of the coupling, one finds the destructive interference regime where the MZMs living in QD1 disappears, giving as a result a trivial zero pole in QD2 and a YSR state with quantized conductance.
@@ Appendix D Superconducting Transport
For computing transport properties in junctions, we start from the general expression for the mean value of the current using Keldysh Green’s functions in Refs. [64, 34]. When considering the current through different SCs, it is convenient to define a Gauge in which the chemical potential difference appears as a time-dependent tunnel coupling, where the SC phase difference evolves as $\phi(t)=\phi_{0}+(2eV/\hbar)t$, cf. Ref. [34]. By a suitable choice of basis, the coupling between the SC probe and a fully polarized Kitaev chain only involves one spin species, equivalent to remove the anomalous part of the SC lead as no multiple Andreev reflections are allowed [69, 33, 34] (i.e., the probe act as a normal lead with a gap). Besides, there is still resonant Andreev transport in the minimal Kitaev chain due to p-wave correlations. Thus, the current through a BCS probe and a spinless p-wave SC in the $2\times 2$ Nambu basis acquires the form,
@@
$G_{RR}(\omega)={}$ $\left[[{\cal G}_{R}]^{-1}-\Sigma(\omega)\right]^{-1},$ (S 18a)
$\Sigma(\omega)={}$ $\Sigma_{LR}^{\dagger}\bar{{\cal G}_{s}}(\omega)\Sigma_{LR}\,,$ (S 18b)
@@
$\bar{{\cal G}_{s}}(\omega)=\operatorname{diag}\left(\frac{-\omega_{-}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega_{-}^{2}}},\frac{-\omega_{+}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega_{+}^{2}}}\right),$ (S 19)
@@
$G^{K}_{RR}(\omega)=G^{\dagger}_{RR}\left\{\Sigma^{K}+[{\cal G}^{\dagger}_{R}]^{-1}{\cal G}^{K}_{R}\ [{\cal G}_{R}]^{-1}\right\}G_{RR}\,,$ (S 20)
@@
${\cal G}^{K}_{R}(\omega)={}$ $(1-2f(\omega))\big{(}{\cal G}^{\dagger}_{R}-{\cal G}_{R}\big{)}\,,$ (S 21a)
$\Sigma^{K}(\omega)={}$ $\Sigma_{LR}^{\dagger}F_{\omega}\big{(}\bar{{\cal G}_{s}}^{\dagger}-\bar{{\cal G}_{s}}\big{)}\Sigma_{LR}\,,$ (S 21b)
@@
$\tilde{\Sigma}^{R/A}(\omega)=\Sigma^{R/A}_{ee}(\omega)\Pi_{e}\,,\quad\tilde{\Sigma}^{K}(\omega)=\Sigma^{K}_{ee}(\omega)\Pi_{e}\,,$ (S 22)
where we have defined the projector $\Pi_{e}=(\sigma_{0}+\sigma_{z})/2$, as only up-spin electrons from the BCS lead are tunnel coupled to the spinless fermions from the minimal Kitaev chain. Finally, we extract the differential conductance from the current using that $G_{S}=\partial I_{S}/\partial V$ in the $T\rightarrow 0$ limit.
The small but finite temperature in the calculation allows for thermally excited quasiparticles from the continuum, inducing a broadening term ($\Gamma_{sp}$) of the Majorana zero pole and enabling single particle current by means of inelastic cotunneling. Furthermore, there is an extra relaxation contribution coming from the non-local coupling to the normal lead. It should be noticed that, for a zero energy pole, both resonant Andreev (i.e., virtual occupation of the state) and single particle processes (i.e., occupation of the state requiring relaxation to empty the state) share the same threshold signal at $e|V|=\Delta$. Notice that, along with the inelastic cotunneling, the hybridization with the SC lead induces a voltage bias dependent broadening ($\Gamma_{e/h}$) of the Majorana zero mode, which may modify the position of the conductance peak related to the threshold. The dominant contribution to the broadening of the subgap state defines two different regimes [33, 72]:
@@
Figure S 5: Non-local effects in SC transport as a function of the coupling to the SC lead. Upper (Lower) panels, conductance for $\varepsilon_{2}/t=0$ ($\varepsilon_{2}/t=1$). Right column represents a close-up image signalling the transition to the ultra-strong coupling. The coupling to the normal lead is set to $\lambda_{n}/t=0.1$ in all panels. Yellow contour marks the universal height value $G_{s}/G_{0}=(4-\pi)$.